Oliver invested $12,000 in an account that accrues compound interest at a rate of 2.5% yearly. What is the minimum number of years Oliver must wait before he has at least double his original investment?
Answer by mananth(16946)

Principal P = 12000
Amount= 24000
n= number of periods 1
compounded (t) 1
Rate = 2.5
n = ln(FV / PV) / ln(1 + r)

n=( ln ( 24000 / 12000 )/ln( 1 + 0.03 )
n=( 0.69 )/ ln 1.03
n=( 0.69 / 0.02 )
n= 28.07
28.07 years